OpenStudy (anonymous):

The value of tolls on the new jersey turnpike is based upon the number of miles traveled. the function can be used to calculate the cost of traveling along the nj turnpike. where x represents the number of miles traveled (rounded to the nearest mile) and t(x) represents the cost of the toll in cents. t(x)= 3x + 15 0 less than or equal to x less than or equal to 95 11x - 725 95< x less than or equal to 112 a.) sketch a graph of t(x) B) find the cost of getting at exit 7A-mile marker 60.9 and getting off at exite 11-mile marker 93.4

OpenStudy (anonymous):

C. find theh cos of traveling the enitr lenght of the NJ turnpike - 112 miles D) approximatley how many miles would you need to travel to have a toll of 2.50

OpenStudy (anonymous):

@master50777 @mathmath333 @perl

OpenStudy (perl):

so plug it into the second part of the piecewise function, since x = 112 miles t(112) = 11*112 - 725

OpenStudy (anonymous):

\[f(x)= \left\{\begin{array}{rcc} 3x+15 & \text{if} & 0\leq x\leq 95\\ 11x-725& \text{if} &95< x < 112 \end{array} \right. \]

OpenStudy (anonymous):

i got 351 for the first one and -602 for the second one

OpenStudy (anonymous):

Are we working on question B?

OpenStudy (anonymous):

check out this widget for a graph the graph is two lines, depending on the domain http://www.wolframalpha.com/widgets/view.jsp?id=5075763dba6ec763f31004272f8aa7fa

OpenStudy (anonymous):

Okay i see the 2 different graphs